How long does it take for 60.0 percent of a sample of radon to decay

1
Expert's answer
2022-01-16T13:09:57-0500

"1-\frac A{A_0}=2^{-\frac tT},"

"t=-T\log_2(1-\frac A{A_0})=5.05^d."